April 3rd 2023

From Hall A Wiki
Jump to: navigation, search
  • Updates
    • ECal
      • No issue with ECAL, two layers went bad because of summing amplifier shower blocks- one module needs to still be repaired
      • data taking looked fine
      • Rpi issue for thresholds from time to time - readback
      • beginning kin 4 not calibrated at beginning
    • Hodoscope
      • issues ?
      • EPICS interface for HV CAEN could be improved
      • add scaler information for hodoscope - FADC or hardware scaler for central bars
      • RF signal : 2 1190 channels - 1 channel looks odd - need to be checked ? Maybe bad connection
    • HCal
      • missing highest energy cluster TDC in HCAL
      • channel missing - can recover TDC information from lower amplitude channel from cluster - but not ideal
      • could still be software
      • is HCAL TDC threshold to low ?
      • need to check Sean for GEn status ?
    • beamline : all looked ok - a few more crashes from sbsvme29 - could check if there are any data issues - will give information